# Witnessing mass-energy equivalence with trapped atom interferometers

### v2

@@ Abstract
Note that while the visibility oscillations are sensitive only to the gravitational correction $\varepsilon_{\text{g}}$, the frequency shift depends both on $\varepsilon_{\text{g}}$ and $\varepsilon_{\text{k}}$. To distinguish between the gravitational and the kinetic effect in fractional frequency shift measurements, one can modulate the separation $\Delta z$ between the upper and the lower trajectory — this will affect the gravitational correction, but not the kinetic one. The probability $P_{g}$ as a function of time is visualized in Fig. 3 for exemplary values of $\omega$, $\omega_{0}$, $\varepsilon_{\text{g}}$, and $\varepsilon_{\text{k}}$.
@@
Figure 3: Oscillations of the (normalized) probability $P_{g}$ for exemplary values $\omega=40\pi$, $\omega_{0}=1.1\omega$, and $\varepsilon_{\text{k}}=0.01$, and two different values of the gravitational correction: $\varepsilon_{\text{g}}=0$ (in gray) and $\varepsilon_{\text{g}}=\varepsilon_{\text{k}}$ (in orange). Note that, apart from visibility modulations, the orange pattern has its frequency shifted against the gray one.
@@
Note that the observation of the frequency shift (6) itself would not be evidence of the quantum clock reading the superposition of proper times since the frequency shift is the same as for the clock being in a statistical mixture of two heights. To reject the latter possibility one can, apart from the total probability $P_{g}$ of measuring the atom in the ground state at the output, extract from the measured data also the difference $D_{g}\equiv P_{g}^{(0)}-P_{g}^{(1)}$ given by (compare with Eq. (1))
$D_{g}=\frac{1}{8}\cos\left(\Delta\phi+\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{u}}}{2}\right),$ (7)
where $\delta_{\text{d}}$ and $\delta_{\text{u}}$ are given in Eq. (4), and $\Delta\phi$ is given by (see Appendix C for derivation)
If the atom is in a superposition of heights, $D_{g}$ will oscillate with time $T_{\text{B}}$. On the other hand, for the atom following a statistical mixture of two heights, the probability difference would be equal to zero. Therefore, extracting from the interferometric output the fractional frequency shift and the oscillations of $D_{g}$ would provide evidence that the atom was reading a superposition of proper times.
It is important to emphasize that the method proposed in this work does not rely on the assumption that ground and excited states follow the same trajectories within the optical lattice. While we showed that within the required order $(1/c^{2})$ of calculating the phases acquired in the interferometer the trajectories associated with different mass-energy states can be approximated as equal, the method allows accounting for mass-energy corrections and resulting phases beyond this approximation.
Finally, let us notice that the interpretation of the experiment as a spectroscopy of a clock at the superposition of heights is possible thanks to including the laser phases and the concrete mechanism that sets up the quantum clock. These were missing in the idealized scenarios in [23, 24] and in the recent work [38] proposing a similar setup using optical tweezers. Therefore, the possibility of measuring the relativistic frequency shift in such interferometric scenarios thus far went unnoticed.
Acknowledgements.
We thank Timothy Le Pers, Navdeep Arya, Germain Tobar, and Jun Ye for helpful discussions and useful comments. The authors acknowledge Knut and Alice Wallenberg foundation through a Wallenberg Academy Fellowship No. 2021.0119. J.F. acknowledges funding from the U.S. Department of Energy, Office of Science, ASCR under Award Number DE-SC0023291 and Australian Research Council Centre of Excellence for Quantum Computation and Communication Technology (No. CE1701000012).
Appendix A Probability calculation
Here we derive the formula (1) for the probability $P_{g}^{(j)}$ of measuring the atom in the ground state in output $j\in\{0,1\}$. We denote by $\Delta\phi$ the phase difference between upper and lower ground state trajectories, and by $\delta_{\text{d}}$ and $\delta_{\text{u}}$ the additional phase acquired by the excited state on the lower and upper trajectory, respectively. The final state of the atom |$\ket{\psi_{\text{final}}}$ can be written in a generic way
|$\ket{\psi_{\text{final}}}=\frac{1}{8}\ket{g}\left(\ket{0}\left[1-\mathrm{e}^{\mathrm{i}\delta_{\text{d}}}+\mathrm{e}^{\mathrm{i}\Delta\phi}\left(1-\mathrm{e}^{\mathrm{i}\delta_{\text{u}}}\right)\right]-\mathrm{i}\ket{1}\left[1-\mathrm{e}^{\mathrm{i}\delta_{\text{d}}}-\mathrm{e}^{\mathrm{i}\Delta\phi}\left(1-\mathrm{e}^{\mathrm{i}\delta_{\text{u}}}\right)\right]\right)+\ldots,$ (9)
where |$\ket{0}$ and |$\ket{1}$ are the states corresponding to two relevant output trajectories, and dots at the end stand for the parts of the state irrelevant to our considerations (lost trajectories and excited state). The factor $1/8=(1/\sqrt{2})^{6}$ in front comes from six laser pulses applied to the atom (four Bragg pulses and two clock pulses). The minus signs in the bracket multiplying |$\ket{0}$ result from the fact that excited state trajectories interacted with a laser two more times than the ground state trajectories (two clock pulses) — each interaction gives rise to the factor $\mathrm{i}\mathrm{e}^{\mathrm{i}\phi_{l}}$ multiplying the corresponding trajectory (here $\phi_{l}$ is the so-called laser phase and we absorb it into phases $\Delta\phi$, $\delta_{\text{d}}$, and $\delta_{\text{u}}$). The factor $-\mathrm{i}$ multiplying |$\ket{1}$ reflects the fact that trajectories leaving the interferometer in output 1 interacted with the odd number of Bragg pulses (the lower trajectory interacted once, the upper one three times), while the ones leaving in output 0 interacted with the even number (two) of Bragg pulses. Finally, the additional minus sign in the bracket multiplying |$\ket{1}$ comes from the fact that, in the case of trajectories leaving in output 1, the upper trajectory interacted with two more Bragg pulses than the lower one.
@@ Appendix A Probability calculation
$\bra{g,j}\ket{\psi_{\text{final}}}\propto\frac{1}{8}\left[1-\mathrm{e}^{\mathrm{i}\delta_{\text{d}}}+(-1)^{j}\mathrm{e}^{\mathrm{i}\Delta\phi}\left(1-\mathrm{e}^{\mathrm{i}\delta_{\text{u}}}\right)\right],$ (10)
where we omitted the common phase factor $-\mathrm{i}$ present in the case of $j=1$. To calculate the probability $P_{g}^{(j)}$ we take the absolute value squared of the above amplitude and get
$P_{g}^{(j)}=\left|\bra{g,j}\ket{\psi_{\text{final}}}\right|^{2}=\frac{1}{16}\left[1-\cos\left(\frac{\delta_{\text{u}}+\delta_{\text{d}}}{2}\right)\cos\left(\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)+(-1)^{j}2\cos\left(\Delta\phi+\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{u}}}{2}\right)\right],$ (11)
@@ Appendix B Fall of the excited-state trajectory
Let us motivate the use of the perturbative method [58] by analyzing the fall of the excited-state trajectory during Bloch oscillations. We assume that the frequencies of the optical lattice lasers are chosen such that the ground state follows a bouncing trajectory at a fixed height (i.e., each interaction with the laser occurs at the same height $z_{\text{B}}$). Then, the excited-state trajectory necessarily falls with each bounce, due to its greater mass. Indeed, since we assume that the interaction with the laser occurs only when the atom reaches downward velocity $v_{\text{B}}=\hbar k/m$, and the momentum transferred to the atom in each interaction is the same for both the ground and the excited state, the excited state will move after each interaction with the upward velocity $v_{\text{B}}-\delta v_{\text{B}}$, where the velocity difference $\delta v_{\text{B}}$ is given by
$\delta v_{\text{B}}=2v_{\text{B}}\frac{\delta m}{m}.$ (12)
@@
$\delta z_{\text{B}}=\frac{v_{\text{B}}\delta v_{\text{B}}}{g}$ (13)
@@
Also, the excited trajectory’s smaller starting velocity means it will reach the downward velocity $v_{\text{B}}$ faster than the ground state. Two subsequent interactions on the excited-state trajectory are separated by a time $\tau_{\text{B}}-\delta\tau_{\text{B}}$, where
$\delta\tau_{\text{B}}=\frac{\delta v_{\text{B}}}{g}.$ (14)
The spatial separation between the ground and the excited state trajectories is maximal when the excited state interacts with the laser. Then, it changes the direction of motion and meets the ground state trajectory at its interaction height $z_{\text{B}}$. Therefore, the trajectories meet repeatedly (at the interaction points of the ground state trajectory), but at each subsequent meeting point the excited-state trajectory has velocity lower by $\delta v_{\text{B}}$ than at the previous meeting point (while the ground state moves there at velocity $v_{\text{B}}$).
[…]
@@ Appendix C Phases
Let us calculate the phases $\Delta\phi$, $\delta_{\text{d}}$, and $\delta_{\text{u}}$ using the perturbative method developed in [58]. Each consists of two contributions: the propagation phase calculated by integrating the free-fall Lagrangian over the trajectory followed by the atom, and the laser phase resulting from atom-light interactions.
@@
Note also that the correction $\delta L^{(2)}$ and the corrections to the laser phases derived in [59] depend only on the trajectory, not on the internal state. Therefore, in the picture in which both the ground and the excited state follow the same trajectory, they will contribute only to the phase $\Delta\phi$, not to $\delta_{\text{d}}$ and $\delta_{\text{u}}$. We are not interested in $1/c^{2}$ corrections to $\Delta\phi$, hence we will omit them in the subsequent analysis and focus only on the corrections coming from $\delta L^{(1)}$.
@@
along the trajectory. More precisely, for the free fall starting at time $t_{\text{A}}$ and ending at $t_{\text{B}}$, the propagation phase is given by
$\begin{split}\phi_{\text{p}}(t_{\text{A}}\to t_{\text{B}})=&\frac{1}{\hbar}\int_{t_{\text{A}}}^{t_{\text{B}}}\mathrm{d}tL_{g}=-\frac{m}{\hbar}\int_{t_{\text{A}}}^{t_{\text{B}}}\mathrm{d}t\left(c^{2}+gz-v^{2}/2\right)\\ =&-\frac{m}{\hbar}\left[\left(c^{2}+gz_{\text{A}}-\frac{1}{2}v_{\text{A}}^{2}\right)(t_{\text{B}}-t_{\text{A}})+v_{\text{A}}g(t_{\text{B}}-t_{\text{A}})^{2}-\frac{1}{3}g^{2}(t_{\text{B}}-t_{\text{A}})^{3}\right],\end{split}$ (20)
@@
$t_{2}-t_{1}=t_{4}-t_{3}=T,\qquad t_{\text{i}}-t_{2}=t_{3}-t_{\text{f}}=T^{\prime},\qquad t_{\text{f}}-t_{\text{i}}=T_{\text{B}}.$ (21)
Let us further denote the quantities corresponding to the lower and upper trajectories by superscripts (d) and (u), respectively, and by $\Delta\phi_{\text{p}}(t_{\text{A}}\to t_{\text{B}})=\phi_{\text{p}}^{\text{(u)}}(t_{\text{A}}\to t_{\text{B}})-\phi_{\text{p}}^{\text{(d)}}(t_{\text{A}}\to t_{\text{B}})$ the propagation phase difference between those two trajectories acquired between $t_{\text{A}}$ and $t_{\text{B}}$. Let us consider one by one the consecutive stages of the interferometer and calculate the corresponding phase difference using (20).
@@
    $\Delta\phi_{\text{p}}(t_{\text{i}}\to t_{\text{f}})=-\frac{m}{\hbar}g\Delta vTT_{\text{B}}.$ (24)
@@
$\Delta\phi_{\text{p}}=\frac{m}{\hbar}\Delta vT\left[v_{1}+\Delta v-v_{3}-g\left(T+2T^{\prime}+T_{\text{B}}\right)\right].$ (27)
@@
On the other hand, $v_{3}$ is the velocity that the atom reaches after time $T^{\prime}$ of a free fall after leaving the optical lattice. Let us notice that the atom does not leave the optical lattice with zero velocity (see Fig. 2), but rather with velocity $-g\tau$. Here $\tau\in\{-\tau_{\text{B}}/2,\tau_{\text{B}}/2\}$ is given by
$T_{\text{B}}=N\tau_{\text{B}}+\tau,$ (29)
where $N=\lfloor T_{\text{B}}/\tau_{\text{B}}+1/2\rfloor$ is the number of interactions between the optical lattice and the atom. Therefore, we have
@@
$\Delta\phi_{\text{p}}=\frac{m}{\hbar}\Delta zg(\tau-T_{\text{B}})=-\frac{m}{\hbar}g\Delta zN\tau_{\text{B}}.$ (31)
@@
Denote by $z_{j}$ with $j\in\{1,2,3,4\}$ the interaction heights with four consecutive Bragg pulses ($z_{1}$ and $z_{3}$ correspond to the upper trajectory, while $z_{2}$ and $z_{4}$ to the lower one), and by $z_{\text{d}}$ and $z_{\text{u}}$ the heights of interactions of the lower and upper trajectories, respectively, within the optical lattice. Note also that the wave vector $k_{\text{Bragg}}$ of the pulses is related to the velocity change $\Delta v$ by
@@
and the wave vector $k_{\text{Bloch}}$ of the optical lattice laser is related to the Bloch period $\tau_{\text{B}}$ by
$2\hbar k_{\text{Bloch}}=mg\tau_{\text{B}},\qquad\Longrightarrow\qquad k_{\text{Bloch}}=\frac{mg\tau_{\text{B}}}{2\hbar}.$ (34)
@@
$\Delta\phi_{\text{l}}=2k_{\text{Bragg}}(z_{1}-z_{2}-z_{3}+z_{4})+2Nk_{\text{Bloch}}(z_{\text{u}}-z_{\text{d}})$ (35)
@@
$z_{2}-z_{1}=v_{1}T-\frac{1}{2}gT^{2},\qquad z_{4}-z_{3}=(v_{3}-\Delta v)T-\frac{1}{2}gT^{2},\qquad z_{\text{u}}-z_{\text{d}}=\Delta z,$ (36)
@@
$\Delta\phi_{\text{l}}=\frac{m}{\hbar}\left[-g\Delta z(T+2T^{\prime}+\tau)+g\Delta zN\tau_{\text{B}}\right].$ (37)
@@
Now, let us calculate the additional phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$ acquired by the excited state on the lower and upper trajectory, respectively. We need to calculate the additional laser phase coming from the clock pulses and the correction to the propagation phase due to the mass-energy equivalence. Starting with the laser phase, we assume that the clock pulses change only the atom’s internal state, but do not affect its motion. This can be achieved by driving two-photon transitions with both photons having the same frequency $\omega/2$ and opposite momenta. The corresponding laser phase is equal to $\mp\omega t$ with the minus (plus) sign corresponding to the two-photon absorption (emission). Note that this phase is independent of the position $z$ of the interaction (thus, it is the same for the lower and upper trajectory). Since the excited state trajectories absorb the photons at $t_{\text{i}}$ and emit at $t_{\text{f}}$, they acquire the laser phase $\omega(t_{\text{f}}-t_{\text{i}})=\omega T_{\text{B}}$.

### v3

@@ Abstract
Note that whereas visibility oscillations are sensitive only to the gravitational correction $\varepsilon_{\rm g}$, the frequency shift depends on both $\varepsilon_{\rm g}$ and $\varepsilon_{\rm k}$. To distinguish the gravitational and kinetic contributions in fractional-frequency-shift measurements, one can modulate the height separation $\Delta z$, which alters $\varepsilon_{\rm g}$ but not $\varepsilon_{\rm k}$. The probability $P_{g}$ as a function of time is shown in Fig. 3 for exemplary values of $\omega$, $\omega_{0}$, $\varepsilon_{\rm g}$, and $\varepsilon_{\rm k}$.
Assuming the experimental parameters as in [20], we would have the velocity $v_{\rm B}\sim 10\,\text{mm/s}$, and the vertical separation of the trajectories $\Delta z\sim 10\,\mu\text{m}$. This would mean that the corrections would be of the order $|\varepsilon_{\rm k}|\sim|\varepsilon_{\rm g}|\sim 10^{-22}$. The fractional frequency shift of that order is very close to what can be currently detected [64, 52, 53, 65]. Indeed, state-of-the-art atomic clocks reach the multiple-measurement precision of the order of $10^{-21}$ [52, 53, 65] (the lowest single-measurement precision in such experiments is $\sim 10^{-19}$ [66]), which is the order of magnitude for the fractional frequency shift in the proposed experiment for separations $\Delta z\sim 100\,\mu\text{m}$, which has already been achieved in [20], at the expense of shorter coherence time.
Figure 3: Oscillations of the (normalized) probability $P_{g}$ for exemplary values $\omega=40\pi$, $\omega_{0}=1.1\omega$, and $\varepsilon_{\rm k}=0.01$, and two different values of the gravitational correction: $\varepsilon_{\rm g}=0$ (in gray) and $\varepsilon_{\rm g}=\varepsilon_{\rm k}$ (in orange). Note that, apart from visibility modulations, the orange pattern has its frequency shifted against the gray one.
On the other hand, for a transition frequency $\omega_{0}\sim 10^{15}\,\mathrm{Hz}$ and a lattice-hold time $T_{\rm B}\sim 1\,\mathrm{s}$ (as in [20]), with $|\varepsilon_{\rm g}|\sim 10^{-22}$, the predicted visibility drop is of order $10^{-15}$. This is far below current experimental sensitivity. Although strontium atoms can support coherent Bloch oscillations for over $100\,\mathrm{s}$ [67], the longest achieved to date is $\sim 1\,\mathrm{s}$. Minute-scale oscillation times have been demonstrated with cesium atoms [21, 22], but their lower $\omega_{0}$ renders the effect even weaker. Extending $T_{\rm B}$ to $100\,\mathrm{s}$ would amplify the visibility oscillations by $10^{4}$, yet they would remain negligible. To achieve an observable visibility drop (of order 0.0110.01-1), one would need a trajectory separation of $\Delta z\sim 10-100\,\mathrm{cm}$. Such separations exceed current capabilities and would likely require two independent lattice lasers.
Note that the observation of the frequency shift (6) itself would not be evidence of the quantum clock reading the superposition of proper times, since the frequency shift is the same as for the clock being in a statistical mixture of two heights. To reject the latter possibility, one can, apart from the total probability $P_{g}$ of measuring the atom in the ground state at the output, extract from the measured data also the difference $D_{g}\equiv P_{g}^{(0)}-P_{g}^{(1)}$ given by (compare with Eq. (1))
$D_{g}=\frac{1}{8}\cos\left(\Delta\phi+\frac{\delta_{\rm u}-\delta_{\rm d}}{2}\right)\sin\left(\frac{\delta_{\rm d}}{2}\right)\sin\left(\frac{\delta_{\rm u}}{2}\right),$ (7)
where $\delta_{\rm d}$ and $\delta_{\rm u}$ are given in Eq. (4), and $\Delta\phi$ reads (see Appendix C for derivation)
If the atom is in a superposition of heights, $D_{g}$ will oscillate with time $T_{\rm B}$. On the other hand, for the atom following a statistical mixture of two heights, the probability difference would be equal to zero. Therefore, extracting from the interferometric output the fractional frequency shift and the oscillations of $D_{g}$ would provide evidence that the atom was reading a superposition of proper times.
It is important to emphasize that the method proposed in this work does not rely on the assumption that the ground and excited states follow the same trajectories within the optical lattice. While we showed that, within the required order $1/c^{2}$ of calculating the phases acquired in the interferometer, the trajectories associated with different mass–energy states can be approximated as equal, the method allows accounting for mass–energy corrections and the resulting phases beyond this approximation. Furthermore, we emphasize that our results do not depend crucially on the simplified Bloch oscillation model we have adopted. Indeed, in Appendix E, we derive analogous results using the usual description of Bloch oscillations [68, 69], in which the atom oscillates under a periodic potential of the optical lattice.
The assumption of zero momentum transfer from the clock pulse is well motivated in special cases, including the usual Lamb-Dicke regime of optical-lattice traps [54, 56, 57, 58], and is exact in the case of two-photon E1-M1 clock transitions [55, 59, 60]. The Lamb-Dicke regime is reached when the atomic motion is limited to a region smaller than the transition wavelength. It was shown [54] that in this case, most transitions occur at the internal transition frequency and do not alter the center-of-mass motion of the atom. However, such behavior is well understood only in the fully quantum picture (with quantized center-of-mass motion) — in our simplified picture with the atom following a semiclassical trajectory, the no-recoil assumption is made as a simplifying yet justified approximation. By contrast, for E1-M1 transitions [55, 59, 60], involving two counter-propagating photons with equal frequencies, the clock transition is recoilless as a rule. In this case, the no-recoil assumption is also well motivated in our simplified model of Bloch oscillations.
Abandoning the assumption of zero momentum transfer naturally leads to recoil shifts [70, 71]. In our model, this implies that the ground and excited states follow different trajectories, and the description becomes significantly more involved. The analysis of this case goes beyond the scope of the present work.
Note that the interpretation of the experiment as spectroscopy of a clock in a superposition of heights is possible only by including the laser phases and the explicit mechanism that establishes the quantum clock. These elements were absent in the idealized scenarios of [23, 24] and in the recent proposal using optical tweezers [38]. Consequently, the possibility of measuring the relativistic frequency shift in such interferometric setups has gone unnoticed until now.
We thank Timothy Le Pers, Navdeep Arya, Germain Tobar, Jun Ye, Kyungtae Kim, and Charles Baynham for helpful discussions and useful comments. The authors acknowledge the Knut and Alice Wallenberg Foundation through a Wallenberg Academy Fellowship No. 2021.0119. J.F. acknowledges funding from the U.S. Department of Energy, Office of Science, ASCR under Award Number DE-SC0023291 and Australian Research Council Centre of Excellence for Quantum Computation and Communication Technology (No. CE1701000012).
Appendix A Probability calculation
@@
[…]
@@
[…]
where we omitted the common phase factor $\mathrm{i}$ present in the case of $j=1$. To calculate the probability $P_{g}^{(j)}$ we take the absolute value squared of the above amplitude and get
[…]
@@ Appendix B Fall of the excited-state trajectory
Let us motivate the use of the perturbative method [62] by analyzing the fall of the excited-state trajectory during Bloch oscillations. We assume that the frequencies of the optical lattice lasers are chosen such that the ground state follows a bouncing trajectory at a fixed height (i.e., each interaction with the laser occurs at the same height $z_{\rm B}$). Then, the excited-state trajectory necessarily falls with each bounce, due to its greater mass. Indeed, since we assume that the interaction with the laser occurs only when the atom reaches downward velocity $v_{\rm B}=\hbar k/m$, and the momentum transferred to the atom in each interaction is the same for both the ground and the excited state, the excited state will move after each interaction with the upward velocity $v_{\rm B}-\delta v_{\rm B}$, where the velocity difference $\delta v_{\rm B}$ is given by
$\delta v_{\rm B}=2v_{\rm B}\frac{\delta m}{m}.$ (12)
@@
$\delta z_{\rm B}=\frac{v_{\rm B}\delta v_{\rm B}}{g}$ (13)
@@
Also, the excited trajectory’s smaller starting velocity means it will reach the downward velocity $v_{\rm B}$ faster than the ground state. Two subsequent interactions on the excited-state trajectory are separated by a time $\tau_{\rm B}-\delta\tau_{\rm B}$, where
$\delta\tau_{\rm B}=\frac{\delta v_{\rm B}}{g}.$ (14)
The spatial separation between the ground and the excited state trajectories is maximal when the excited state interacts with the laser. Then, it changes the direction of motion and meets the ground state trajectory at its interaction height $z_{\rm B}$. Therefore, the trajectories meet repeatedly (at the interaction points of the ground state trajectory), but at each subsequent meeting point the excited-state trajectory has velocity lower by $\delta v_{\rm B}$ than at the previous meeting point (while the ground state moves there at velocity $v_{\rm B}$).
The largest separation between the $n$th and $(n+1)$th interaction on the ground state trajectory occurs at the moment of $(n+1)$th interaction on the excited-state trajectory and is equal to $2n\delta z_{\rm B}$. For the experimental parameters from [20] (strontium atoms, optical lattice operating at $532\;\text{nm}$, around 500 oscillations) the maximal separation between the ground and excited state trajectories would be around $10^{-13}\;\text{m}$. This should be compared with the thermal wavelength of the atoms given by
@@ Appendix C Phases
Let us calculate the phases $\Delta\phi$, $\delta_{\rm d}$, and $\delta_{\rm u}$ using the perturbative method developed in [62]. Each consists of two contributions: the propagation phase calculated by integrating the free-fall Lagrangian over the trajectory followed by the atom, and the laser phase resulting from atom-light interactions.
@@
Note also that the correction $\delta L^{(2)}$ and the corrections to the laser phases derived in [63] depend only on the trajectory, not on the internal state. Therefore, in the picture in which both the ground and the excited state follow the same trajectory, they will contribute only to the phase $\Delta\phi$, not to $\delta_{\rm d}$ and $\delta_{\rm u}$. We are not interested in $1/c^{2}$ corrections to $\Delta\phi$, hence we will omit them in the subsequent analysis and focus only on the corrections coming from $\delta L^{(1)}$.
@@
along the trajectory. More precisely, for the free fall starting at time $t_{\text{A}}$ and ending at $t_{\rm B}$, the propagation phase is given by
$\begin{split}\phi_{\text{p}}(t_{\text{A}}\to t_{\rm B})=&\frac{1}{\hbar}\int_{t_{\text{A}}}^{t_{\rm B}}\mathrm{d}tL_{g}=-\frac{m}{\hbar}\int_{t_{\text{A}}}^{t_{\rm B}}\mathrm{d}t\left(c^{2}+gz-v^{2}/2\right)\\ =&-\frac{m}{\hbar}\left[\left(c^{2}+gz_{\text{A}}-\frac{1}{2}v_{\text{A}}^{2}\right)(t_{\rm B}-t_{\text{A}})+v_{\text{A}}g(t_{\rm B}-t_{\text{A}})^{2}-\frac{1}{3}g^{2}(t_{\rm B}-t_{\text{A}})^{3}\right],\end{split}$ (20)
@@
$t_{2}-t_{1}=t_{4}-t_{3}=T,\qquad t_{\text{i}}-t_{2}=t_{3}-t_{\text{f}}=T^{\prime},\qquad t_{\text{f}}-t_{\text{i}}=T_{\rm B}.$ (21)
Let us further denote the quantities corresponding to the lower and upper trajectories by superscripts (d) and (u), respectively, and by $\Delta\phi_{\text{p}}(t_{\text{A}}\to t_{\rm B})=\phi_{\text{p}}^{\text{(u)}}(t_{\text{A}}\to t_{\rm B})-\phi_{\text{p}}^{\text{(d)}}(t_{\text{A}}\to t_{\rm B})$ the propagation phase difference between those two trajectories acquired between $t_{\text{A}}$ and $t_{\rm B}$. Let us consider one by one the consecutive stages of the interferometer and calculate the corresponding phase difference using (20).
@@
    $\Delta\phi_{\text{p}}(t_{\text{i}}\to t_{\text{f}})=-\frac{m}{\hbar}g\Delta vTT_{\rm B}.$ (24)
@@
$\Delta\phi_{\text{p}}=\frac{m}{\hbar}\Delta vT\left[v_{1}+\Delta v-v_{3}-g\left(T+2T^{\prime}+T_{\rm B}\right)\right].$ (27)
@@
On the other hand, $v_{3}$ is the velocity that the atom reaches after time $T^{\prime}$ of a free fall after leaving the optical lattice. Let us notice that the atom does not leave the optical lattice with zero velocity (see Fig. 2), but rather with velocity $g\tau$. Here $\tau\in\{-\tau_{\rm B}/2,\tau_{\rm B}/2\}$ is given by
$T_{\rm B}=N\tau_{\rm B}+\tau,$ (29)
where $N=\lfloor T_{\rm B}/\tau_{\rm B}+1/2\rfloor$ is the number of interactions between the optical lattice and the atom. Therefore, we have
@@
$\Delta\phi_{\text{p}}=\frac{m}{\hbar}\Delta zg(\tau-T_{\rm B})=-\frac{m}{\hbar}g\Delta zN\tau_{\rm B}.$ (31)
@@
Denote by $z_{j}$ with $j\in\{1,2,3,4\}$ the interaction heights with four consecutive Bragg pulses ($z_{1}$ and $z_{3}$ correspond to the upper trajectory, while $z_{2}$ and $z_{4}$ to the lower one), and by $z_{\rm d}$ and $z_{\rm u}$ the heights of interactions of the lower and upper trajectories, respectively, within the optical lattice. Note also that the wave vector $k_{\text{Bragg}}$ of the pulses is related to the velocity change $\Delta v$ by
@@
and the wave vector $k_{\text{Bloch}}$ of the optical lattice laser is related to the Bloch period $\tau_{\rm B}$ by
$2\hbar k_{\text{Bloch}}=mg\tau_{\rm B},\qquad\Longrightarrow\qquad k_{\text{Bloch}}=\frac{mg\tau_{\rm B}}{2\hbar}.$ (34)
@@
$\Delta\phi_{\text{l}}=2k_{\text{Bragg}}(z_{1}-z_{2}-z_{3}+z_{4})+2Nk_{\text{Bloch}}(z_{\rm u}-z_{\rm d})$ (35)
@@
$z_{2}-z_{1}=v_{1}T-\frac{1}{2}gT^{2},\qquad z_{4}-z_{3}=(v_{3}-\Delta v)T-\frac{1}{2}gT^{2},\qquad z_{\rm u}-z_{\rm d}=\Delta z,$ (36)
@@
$\Delta\phi_{\text{l}}=\frac{m}{\hbar}\left[-g\Delta z(T+2T^{\prime}+\tau)+g\Delta zN\tau_{\rm B}\right].$ (37)
@@
Now, let us calculate the additional phases $\delta_{\rm d}$ and $\delta_{\rm u}$ acquired by the excited state on the lower and upper trajectory, respectively. We need to calculate the additional laser phase coming from the clock pulses and the correction to the propagation phase due to the mass-energy equivalence. Starting with the laser phase, we assume that the clock pulses change only the atom’s internal state, but do not affect its motion. This can be achieved by driving two-photon transitions with both photons having the same frequency $\omega/2$ and opposite momenta. The corresponding laser phase is equal to $\mp\omega t$ with the minus (plus) sign corresponding to the two-photon absorption (emission). Note that this phase is independent of the position $z$ of the interaction (thus, it is the same for the lower and upper trajectory). Since the excited state trajectories absorb the photons at $t_{\text{i}}$ and emit at $t_{\text{f}}$, they acquire the laser phase $\omega(t_{\text{f}}-t_{\text{i}})=\omega T_{\rm B}$.